OpenStudy (anonymous):

prove mathematical induction (sum 2^0 + 2^1 + ...+ 2^n is 2^n+1 - 1 for all n>=1 how can i prove???

OpenStudy (anonymous):

prove for case \(n=1\) is easy enough right?

OpenStudy (anonymous):

now we assume it is true if \(n=k\) , that is, we assume \[2^2+2^1+2^2+...+2^k=2^{k+1}-1\] and using that fact we want to show it is true for \(n=k+1\)

OpenStudy (anonymous):

what does the \(k+1\) statement say? it says \(2^0+2^1+2^2+...+2^k+2^{k+1}=2^{k+2}-1\) so that is what we need to work towards

OpenStudy (anonymous):

ok i understand it but how explain step wise ??

OpenStudy (anonymous):

\[2^0+2^1+2^2+...+2^k+2^{k+1}=2^{k+2}-1\] \[\overbrace{2^0+2^1+2^2+...+2^k}+2^{k+1}=2^{k+2}-1\] by induction we can replace that part in indicated by the formula we get to assume

OpenStudy (anonymous):

ok typo there sorry

OpenStudy (anonymous):

u r great dude.....thanks

OpenStudy (anonymous):

\[\overbrace{2^0+2^1+2^2+...+2^k}+2^{k+1}=\overbrace{2^{k+1}-1}+2^{k+1}\]

OpenStudy (anonymous):

equating the stuff under the braces is the induction hypothesis the rest is algebra or arithmetic. make the right hand side look like \(2^{k+2}-1\)

OpenStudy (anonymous):

it always works this way with induction and summation formulas. just break off the last term to use induction on the rest up until the last term, then do some algebra

OpenStudy (anonymous):

oh and yw, hope all steps are clear